# An Empirical Analysis on the Use and Reporting of National Security Letters

### v3

@@ 3 Data Collection and Trends
Despite various reports published by government agencies (mandated by law) and companies that aim to increase the transparency and auditability of NSL usage, interpreting and auditing these reports can be a tiresome process. Concretely, there is significant manual effort required to identify, collect, clean, normalize and interpret the range of public NSL reports. There exists no single collection of NSL information; finding reports from various entities is not always straightforward and can be time-consuming. Additionally, there is no universal format that all reports follow. They can be formatted differently depending on the agency that published them, or when they were published. Company transparency reports are even less structured, as they are reported voluntarily by different companies. Moreover, this is not merely a matter of format, but the kind of data published may change over time (even within the publication portfolio of a single organization), the semantics of its interpretation (e.g., whether or not one reports that precisely 0 NSLs were received), and the granularity of aggregation. Moreover, data may overlap, such as when NSLs with the same number are served to multiple companies. [^4] For instance, NSL-19-483160 is served to both Google and Apple.
@@
One contribution of this paper is the first comprehensive collection of NSL statistics, transparency reports, and published NSL letters scattered across the Internet. In addition, using automated scripts and manual cleaning, we extract all useful information from the NSL reports into a normalized format, enabling several analyses detailed in later sections. We make our data set and scripts public. [^5] We publish our data set here: https://github.com/ucsdsysnet/nsl-empirical-analysis.
@@
Figure 1 provides an overview of the time ranges [^6] These ranges do not necessarily correspond to amendments. For instance, companies only received permission to publish statistics on the number of transparency reports they received after an agreement with the DOJ in 2014 [15]. However, some companies retroactively published their NSL statistics back to 2009 after 2014. and the type of information that we were able to collect from different sources:
@@
  • type 1 (cf. Section 3.1): Multiple government entities publish statistics about NSL usage (type 1 data): The OIG reports from 2006 [36], 2007 [38], and 2014 [37] contain the sum of all types of ROIs for 2003–2011, and US and non-US ROIs for 2003–2009. FISA reports [35] contain NSL requests and ROIs for US targets since 2005, and data for non-US and subscriber information requests since 2015.
@@
  • type 3 (cf. Section 3.3): After the gag order was lifted, companies may publish redacted NSL letters that they received. Albeit personally identifiable information is redacted, the letters still contain useful metadata including file numbers and issuance dates. In most cases, companies document when they published an NSL.
Figure 1: Overview of data available from three types of sources: government statistics (type 1), transparency reports (type 2), and published NSLs (type 3)
Broadly speaking, the reports we collect come from three sources. First, government reports such as statistics published by the FBI. Second, transparency reports from companies that contain the number of NSL requests they received. Third, redacted NSLs published by companies after their gag order was lifted. Below, we detail how we collect and clean data from each source and observe general trends.
3.1 Government Reports
NSL statistics collected from various government reports [34, 35, 36, 38, 37] include the number of NSL requests for US and non-US persons, as well as subscriber information for any person, including both US and non-US persons. Figure 2 depicts this data, enumerating the number of targeted individuals for every type of NSL request. A single NSL can contain multiple requests for information, or ROIs, as long as they are in the context of a single investigation. As well, we note that the FBI may serve multiple NSLs for the same person under different statutes and to different entities (e.g. to collect information on telephone numbers, email addresses, and financial records) and thus the number of NSL requests can be greater than the number of targets.
@@ 3.1 Government Reports
The FISA reports [35, 3] started to mandate reporting NSL requests for US persons in 2003. However, the format of the reported data changed multiple times since then due to new regulations and policies. In 2005, section 128 of the USA PATRIOT Improvement and Reauthorization Act added the requirement to report the number of ROIs for US persons made with NSLs [14]. Starting in 2015, the USA FREEDOM Act [19] additionally required the reporting of NSL requests for non-US persons and NSL requests for subscriber information. The ASTR [34] published by the ODNI reports the total number of issued NSLs and ROIs.
@@
For 2003–2009, we found reports of the OIG [38, 37] to contain statistics (that were redacted in an early version but disclosed in the 2014 revision [37]) from reports of the FBI to Congress under FISA on the number of NSL requests for non-US persons (shown in yellow in Figure 2). The OIG reports also include the total number of NSL requests until 2011, from which we can infer the number of non-US ROIs for 2003–2009 and the combined number of non-US ROIs and subscriber information requests for 2010 and 2011. Unfortunately, the OIG reports do not contain any additional information about non-US NSL targets. Thus, for the period from 2012 to 2014, we only know statistics about NSL requests and targets pertaining to US persons due to the lack of declassified information.
Figure 2: Number of requests and targets for different NSL types from 2003 to 2022. For every year, the left bar shows the type of targets in NSL requests, starting from 2005: US persons, non-US persons, or letters for subscriber information (for which the target’s nationality is unknown a priori). The right bar is aggregated from various sources and shows the types of Requests of Information (ROI; an NSL can contain multiple of them), split into the same categories.
@@
Furthermore, Figure 2 indicates that the number of NSL requests for US persons grew steadily until 2010, and then eventually decreased back to the level of 2003. Partial data from OIG reports [38, 37] suggests that NSL requests for non-US persons were in the minority before 2010, but became more common between 2010 and 2015 and still remain popular. Additional data from the FISA reports shows that after 2015, the US-targeted NSL requests represent a smaller portion of all requests, although a significant portion of requests are made for subscriber information of any person. [^7] It is tempting to relate fluctuations in NSL requests to global political events. For instance, the peak in 2019 for non-US ROIs may be due to increased investigations into digital attacks to perform economic espionage, which dominated the FBI’s counterintelligence program in 2019 according to FBI director Christopher Wray [46]. However, it is impossible to corroborate such a hypothesis using only the data available here.
3.2 Transparency Reports
@@
Transparency reports provide the perspectives of individual companies on NSL issuance. In the United States, there is neither a legal requirement nor a standardized format for publishing transparency reports. Moreover, the government restricts the reporting of the number of NSL requests to bands of 250, 500, or 1000 as defined by 50 U.S.C. § 1874 [18]. These restrictions and lack of structure cause the number, frequency, content, and accessibility of transparency reports to vary wildly by company.
@@
We collect transparency reports published by 55 US-based private companies (listed in Appendix A.1), of which 41 provided data related to NSLs. While some of the companies have all of their transparency reports easily archived and downloadable as CSV or PDF files, other reports are hidden in blog posts or support forum answers, posted as low-quality images, or have broken links. In addition, we had to exclude the reports of seventeen companies as they only released aggregated numbers for all national security requests they received, mixing NSLs with other FISA requests (e.g., for electronic surveillance). Clearly, our dataset is not comprehensive and only provides lower bounds on the NSL usage. It is possible that we missed some transparency reports in our search, and the data of companies that choose not to disclose this information is absent.
@@
Figure 3: The highlighted area shows the accumulated transparency report ranges, computed as the sum of lower and upper limits reported by some companies. We compare them to the line of government-reported number of issued NSLs from the ASTR.
@@
Despite the incomplete data set and that companies report the number of NSLs in coarse bands, we can still compare the sum of lower respective upper bounds with the total number of issued NSLs from ASTR as shown in Figure 3. We find that the number of NSL requests from ASTR exceeds the cumulative upper bound from transparency reports, indicating that we are indeed missing a significant fraction of company reports for NSLs. Although the company transparency records are not a comprehensive data source, we observe that after 2013, the mean of the bands strongly correlates with the contemporaneous number of issued NSLs (with a Pearson coefficient of 0.77). This indicates that the two data sources are, at the very least, not inconsistent. However, more fine-grained data or more standardized transparency reports are needed to support stronger conclusions.
@@
The growing distance between cumulative lower and upper bands in Figure 3 from 2009 to 2013 is due to the growing number of published transparency reports. Initially, only Google and Twitter published reports. There was a sharp spike in the number of companies publishing reports around 2013 and 2014, presumably in response to the Snowden revelations drawing increased attention to transparency. Some companies stopped issuing reports since then, but the decrease in reports in 2022, which is reflected in the decreased width of the band, is likely an artifact of delayed report publishing by companies.
@@
Diving into the reported ranges of NSLs themselves, we find that 13 out of 41 reporting companies explicitly stated that they had never received any NSLs. The vast majority of companies report the lowest band for NSL requests (i.e. 0-249, 0-499, or 0-999). Telecommunication companies, such as AT&T, T-Mobile, and Verizon, received the most NSLs, with up to 1000-2000 requests each year. Apple and Google, both large producers of phones and mobile operating systems, were the only other companies reporting a higher range of NSL requests. We surmise that telecommunications and related companies may lead these statistics because they collect valuable metadata for investigations.888 In comparison, companies like Adobe are less likely to be utilized by malicious actors in a way that produces useful metadata for investigations.
3.3 Company NSLs
@@
The third and last data source are NSLs themselves, voluntarily published by their recipients [^9] Two avenues for publishing NSLs are the reciprocal notice procedure and the termination procedure. after the nondisclosure requirement (colloquially “gag order”) has been lifted. These letters request specific account information (called ROIs) and have a gag order attached, restricting the receiving party from publishing the orders or even discussing their existence. The most sensitive parts of the letters are redacted, including user identities and confidential information such as Social Security Numbers. However, unredacted information includes the types of information requested (e.g. email metadata or credit records) and types of user identifiers (e.g. email addresses, account numbers, addresses, or names). Additionally, we found file numbers and publication and issuance dates (which are often not redacted) to be useful metadata, as well as that the number of redacted lines indicates the ROI volume.
@@
Google and Apple are the largest publishers of NSL letters in our data set with 272 and 45 letters, respectively. Appendix A.2 lists the full composition of our data set. We only see a fraction of all existing letters. Between Jan 1, 2015 and Dec 31, 2020101010We pick these years such that they are after the gag order was weakened but not too recent to avoid too much skew because of non-disclosure orders that are still preventing companies from publishing the letters. , companies published only 0.3% of all issued NSLs. We discuss interesting trends that can still be observed from the available data in the remainder of this section.
@@
We use the number of days between the issuance and publication dates as an estimator for the duration of the gag order. The company-internal administrative delay between lifting a gag order and publishing the letter may add some noise to this metric.
@@
The scatter plot in Figure 4 shows for all public NSLs, how many years after their issuance that companies publish them. We observe that, except for two NSLs, all letters were published after June 2015, when the USA FREEDOM Act [19] amended NSL reporting and nondisclosure regulations, limiting gag orders to the duration for which publication would interfere with national security interests or ongoing investigations, and simplifying the disclosure process. This supports, that in the absence of public scrutiny, the gag orders before 2015 were easier to enforce and very rarely lifted. After Congress relaxed the nondisclosure requirements, publishing NSLs became possible for companies. It appears that few companies retroactively published NSLs that they received before these legal reforms.
@@
Figure 4: This graph shows how many years it took to publish NSLs after they were issued. Most letters were only published after the USA FREEDOM Act added procedures to challenge gag orders. The gag order time appears to decrease. However, the distribution of the unpublished 99.7% of the NSLs is unknown. Moreover, the measured gag order time appears to decrease due to a bias, as letters from recent years with longer times have not been published yet. For comparison, we plot a black line to show a trend for a simulated scenario where gag order time is sampled from the uniform distribution with the constant mean value of 5 years.
@@
However, two important caveats need to be mentioned in connection with Figure 4. First, we only see 0.3% of all NSLs, and it is unknown whether they are a representative sample for the distribution of the remaining 99.7% of NSLs. Second, data since 2013 is increasingly skewed towards shorter gag order times, as any letter that is in the “future dates” area will only be published in the future. For instance, letters issued in 2022 with a gag order time of four years will only be published in 2026. We characterize the bias and provide a toy model to estimate the true, unbiased mean gag order duration time in Appendix B.
An additional challenge in interpreting NSL data is the unknown bias stemming from possibly selective publishing, due to which gag order the government decides to lift and which letters companies voluntarily publish. Companies have no obligation to publish NSLs, and we have little insight into their internal reasoning for whether or not to disclose NSLs: there might be a selection bias in the volume and type of NSLs they publish.
4 Data Inconsistencies

### v4

@@ 4. Data Collection and Trends
Despite that the published data should improve transparency and auditability of NSL usage, interpreting and auditing these reports is a tiresome process. Significant manual effort is required to identify, collect, clean, normalize and interpret the range of public NSL reports. There exists no single collection of NSL information, and finding reports from various entities is not always straightforward or efficient. Additionally, there is no universal format that all reports follow. Company transparency reports are even less structured, as they are reported voluntarily and without any standards defining format, frequency of publication, or required information. Moreover, data may overlap, such as when NSLs with the same number are served to multiple companies. [^3] For instance, NSL-19-483160 is served to both Google and Apple.
@@
One contribution of this paper is the first comprehensive collection of NSL statistics, transparency reports, and published NSL letters scattered across the Internet. Additionally, using automated scripts and manual cleaning, we extract all useful information from the NSL reports into a normalized and computer-friendly format, enabling several analyses detailed in later sections. We make our data and scripts public. [^4] We publish our data set here: https://github.com/ucsdsysnet/nsl-empirical-analysis.
@@
Figure 1 shows the time ranges [^5] These ranges do not necessarily correspond to amendments. For instance, companies only received permission to publish statistics on the number of transparency reports they received after an agreement with the DOJ in 2014 (dojagreement:2014, ). However, after 2014, some companies retroactively published their NSL statistics back to 2009. and information that we were able to collect from different sources:
@@
  • Type 1 (cf. Section 4.1): Multiple government entities publish statistics about NSL usage (type 1 data): The OIG reports from 2006 (ig-report-2006, ), 2007 (ig-report-2007, ), and 2014 (ig-report-2014, ) contain the sum of all types of ROIs for 2003–2011, and US and non-US ROIs for 2003–2009. FISA reports (fisa-reports, ) contain NSL requests and ROIs for US targets since 2005, and data for non-US and subscriber information requests since 2015.
@@
  • Type 3 (cf. Section 4.3): After associated gag orders are lifted, companies may publish redacted NSL letters that they received. While personally identifiable information is redacted, the letters still contain useful metadata including file numbers and issuance dates. In most cases, companies document when they published an NSL.
Below, we detail how we collect and clean data from each source and observe general trends.
4.1. Government Reports
NSL statistics collected from various government reports (astr, ; fisa-reports, ; ig-report-2006, ; ig-report-2007, ; ig-report-2014, ) include the number of NSL requests for US and non-US persons, as well as subscriber information for any person. Figure 2 depicts this data. The number of targeted individuals for every type of NSL request is enumerated in Figure 2(a). Figure 2(b) shows the number of ROIs, which exceeds the number of targets as one person can have multiple NSLs issued under different statutes, each containing multiple ROIs. While the published information does not allow us to associate NSL letters with their respective number of ROIs, we can give overall statistics. Since 2015, there were on average 3.8 ROIs per target, with a standard deviation of 0.7. The maximum was in 2015, at 5.3 ROIs per target.
(a) Number of targets.
(b) Number of ROIs.
Figure 2. Reported information about NSL requests for US persons, non-US persons, and subscriber information (abbreviated as “sub. info”) in case the target’s nationality is unknown at the time of request. A single NSL letter can request data for multiple ROIs (e.g., email addresses) of the same target.
[…]
[ Two bar plots show the number of targets as well as the number of ROIs between 2003 and 2022.] Reported information about NSL requests for US persons, non-US persons, and subscriber information (abbreviated as “sub. info”) in case the target’s nationality is unknown at the time of request. A single NSL letter can request data for multiple ROIs (e.g., email addresses) of the same target.
@@ 4.1. Government Reports
The FISA reports (fisa-reports, ; law:fisaReport, ) started to mandate reporting NSL requests for US persons in 2003. Since then, the format of the reported data has changed multiple times due to new regulations and policies. In 2005, section 128 of the USA PATRIOT Improvement and Reauthorization Act added the requirement to report the number of ROIs for US persons made with NSLs (law:PL-109-177-PATRIOT-2005-section-128, ). Starting in 2015, the USA FREEDOM Act (law:PL-114-23-FREEDOM, ) additionally required the reporting of NSL requests for non-US persons and NSL requests for subscriber information. The ASTR (astr, ) published by the ODNI reports the total number of issued NSLs and ROIs.
@@
For 2003–2009, we found that reports of the OIG (ig-report-2007, ; ig-report-2014, ) contained statistics (that were redacted in an early version but disclosed in the 2014 revision (ig-report-2014, )) from reports of the FBI to Congress under FISA on the number of ROIs in NSLs for US and non-US persons. The OIG reports also include the total number of NSL requests until 2011, from which we can infer the number of subscriber information ROIs for 2003–2009 and the combined number of non-US ROIs and subscriber information requests for 2010 and 2011. Unfortunately, the OIG reports do not contain any additional information about non-US or subscriber information NSLs. Thus, for the period from 2012 to 2014, we only know statistics about ROIs pertaining to US persons due to the lack of declassified information (see Figure 2(b)). Similarly, we do not have any information on the targets of NSLs for non-US and subscriber information before 2015 as shown in (see Figure 2(a)).
@@
Figure 2 indicates that the number of NSL requests for US persons grew steadily until 2010, and then eventually decreased back to the level of 2003. Partial data from OIG reports (ig-report-2007, ; ig-report-2014, ) suggests that NSL requests for non-US persons were in the minority before 2010, but became more common between 2010 and 2015 and still remain popular. Additional data from the FISA reports shows that after 2015, the US-targeted NSL requests represent a smaller portion of all requests, although a significant portion of requests are made for subscriber information of any person. [^6] It is tempting to relate fluctuations in NSL requests to global political events. For instance, the peak in 2019 for non-US ROIs may be due to increased investigations into digital attacks to perform economic espionage, which dominated the FBI’s counterintelligence program in 2019 according to FBI director Christopher Wray (counterintelligence-2019, ). However, it is impossible to corroborate such a hypothesis using only the data available here.
4.2. Transparency Reports
@@
Transparency reports provide the perspectives of individual companies on NSL issuance. In the United States, there is neither a legal requirement nor a standardized format for publishing transparency reports. Moreover, the government restricts the reporting of the number of NSL requests to bands of 250, 500, or 1000 as defined by 50 U.S.C. § 1874 (law:USC-50-1874-NSL-ranges, ). These restrictions and lack of structure cause the number, frequency, content, and accessibility of transparency reports to vary wildly by company and year.
We collect transparency reports published by 55 US-based private companies (listed in Appendix B.1), of which 41 provided data related to NSLs. These companies were selected from Access Now’s Transparency Reporting Index777www.accessnow.org/campaign/transparency-reporting-index, which to our knowledge is the only public database of companies that publish transparency reports. We removed 33 companies listed in the Index from consideration as they were not US-based companies, and therefore did not publish any information related to US national security inquiries. Given the lack of centralized data about transparency reports, our findings necessarily represent a lower bound for the number of NSL requests issued to US companies.
@@
While some of the companies have all of their transparency reports easily archived and downloadable as CSV or PDF files, other reports are hidden in blog posts or support forum answers, posted as low-quality images, or have broken links. In addition, we had to exclude the reports of 17 companies as they only released aggregated numbers for all national security requests they received, mixing NSLs with other FISA requests (e.g., for electronic surveillance). Our dataset is therefore not comprehensive and only provides lower bounds on NSL usage.
@@
Figure 3. The highlighted area shows the cumulative transparency report ranges, the sum of limits reported by some companies. We compare them to the number of issued NSLs reported by the ASTR.
[…]
[This visualization shows the cumulative transparency report ranges, the sum of limits reported by some companies. We compare them to the number of issued NSLs reported by the ASTR.] The plot shows that the number of issued NSLs is most of the times higher than the cumulative upper bound of the limits, showing that there is missing data. However, the inclines and declines seem consistent with the changes in the cumulative bounds.
@@
Despite the incomplete data set and coarse reporting bands, we can still compare the respective sums of lower and upper bounds with the total number of issued NSLs from ASTR as shown in Figure 3. We find that the number of NSL requests from ASTR exceeds the cumulative upper bound from transparency reports, indicating that we are indeed missing a significant fraction of company reports for NSLs. Although the company transparency records are not a comprehensive data source, we observe that after 2013, the mean of the bands strongly correlates with the contemporaneous number of issued NSLs (with a Pearson coefficient of 0.77). This indicates that the two data sources are, at the very least, not inconsistent. However, more fine-grained data is needed to support stronger conclusions.
@@
The growing distance between cumulative lower and upper bands in Figure 3 from 2009 to 2013 is due to the growing number of published transparency reports. Initially, only Google and Twitter published reports. There was a sharp spike in the number of companies publishing reports around 2013 and 2014, presumably in response to the Snowden revelations drawing increased attention to transparency. Some companies have stopped issuing reports since then, but the decrease in reports in 2022, which is reflected in the decreased width of the band, is likely an artifact of delayed transparency report publishing by companies.
@@
Diving into the reported ranges of NSLs themselves, we find that 13 out of 41 reporting companies explicitly stated that they had never received any NSLs. The vast majority of companies report the lowest band for NSL requests (i.e., 0-249, 0-499, or 0-999). Telecommunication companies, such as AT&T, T-Mobile, and Verizon, received the most NSLs, with up to 1000-2000 requests each year. Apple and Google, both large producers of phones and mobile operating systems, were the only other companies reporting a higher range of NSL requests. We surmise that telecommunications and related companies may lead these statistics because they collect valuable metadata for investigations.888 In comparison, companies like Adobe maybe be less likely to be utilized by malicious actors.
Due to legal restrictions regarding the discussion of NSLs by recipient companies, we were not able to reach out to any companies to ask more detailed questions about specific NSLs or general patterns regarding NSLs that they have observed. We were, however, able to investigate the subset of NSLs that have been published by some companies once the NSLs are no longer under nondisclosure restrictions. Analyzing these NSLs is the focus of Section 4.3.
4.3. Company NSLs
@@
The third and final data source are NSLs themselves, voluntarily published by their recipients [^9] Two avenues for publishing NSLs are the reciprocal notice procedure and the termination procedure. after the nondisclosure requirement (colloquially “gag order”) has been lifted. These letters contain specific ROIs and have a gag order attached, restricting the receiving party from publishing the NSLs or even discussing their existence. The most sensitive parts of the letters are redacted, including user identities and confidential information such as Social Security Numbers. However, unredacted information includes the types of information requested (e.g., email metadata or credit records) and types of ROIs (i.e., user identifiers such as email addresses, account numbers, addresses, or names). Additionally, we found file numbers and the issuance date (which are often not redacted) to be useful metadata, in addition to the number of redacted lines that indicate the ROI volume. Finally, most companies report the publication date on which they released the redacted NSL.
@@
Google and Apple are the largest publishers of NSL letters in our data set with 272 and 45 letters, respectively. Appendix B.2 lists the full composition of our data set. We only see a fraction of all existing letters: Between Jan 1, 2015 and Dec 31, 2020101010We pick these years as they are after the gag order was weakened but not too recent include currently-active gag orders. , companies published only 0.3% of all issued NSLs.
@@
We use the number of days between the issuance and publication dates as an estimator for the duration of the gag order. The internal administrative delay between lifting a gag order and publishing the letter may add some noise to this metric.
@@
The scatter plot in Figure 4 shows how many years after an NSL’s issuance it is published. We observe that, except for two NSLs, all letters were published after June 2015, when the USA FREEDOM Act (law:PL-114-23-FREEDOM, ) amended NSL reporting and nondisclosure regulations, limited gag orders, and simplified the disclosure process. This supports the hypothesis that in the absence of public scrutiny, the gag orders before 2015 were easier to enforce and very rarely lifted. After Congress relaxed the nondisclosure requirements, publishing NSLs became possible for companies. It appears that few companies retroactively published NSLs that they received before these legal reforms.
@@
Figure 4. This graph shows how many years it took to publish NSLs after they were issued. Most letters were only published after the USA FREEDOM Act added procedures to challenge gag orders. The gag order time appears to decrease. However, the distribution of the unpublished 99.7% of the NSLs is unknown. Moreover, the measured gag order time appears to decrease due to a bias, as letters from recent years with longer times have not been published yet. For comparison, we plot a black line to show a trend for a simulated scenario where gag order time is sampled from the uniform distribution with the constant mean value of 5 years.
[…]
[A graph plotting NSLs with their issuance date on the x axis and the number of years until they were published on the y axis.] This graph shows how many years it took to publish NSLs after they were issued. Most letters were only published after the USA FREEDOM Act added procedures to challenge gag orders. The gag order time appears to decrease. However, the distribution of the unpublished 99.7% of the NSLs is unknown. Moreover, the measured gag order time appears to decrease due to a bias, as letters from recent years with longer times have not been published yet. For comparison, we plot a black line to show a trend for a simulated scenario where gag order time is sampled from the uniform distribution with the constant mean value of 5 years.
@@
There are two important caveats related to Figure 4. First, we only see 0.3% of all NSLs, and it is unknown whether they are a representative sample for the distribution of the remaining 99.7% of NSLs. Second, data since 2013 is increasingly skewed towards shorter gag order times, as any letter that is in the “future dates” area will only be published in the future. For instance, letters issued in 2022 with a gag order time of four years will only be published in 2026. We characterize this bias with the simulated gag order time (black line) in Figure 4 compared to the polynomial regression of observed gag order times (red line). Furthermore, we provide a toy model to estimate the true, unbiased mean gag order duration time in Appendix C.
5. Data Inconsistencies